# Two-Stage Pretraining for Molecular Property Prediction in the Wild

### v1

@@ 3.2 Stage 2: Auxiliary Property Prediction
We further improve the generalization capability of the primary encoder by incorporating auxiliary property prediction in the second pretraining stage. This approach is inspired by multitask learning (Caruana, 1997), where a model is trained to solve both the primary task and related auxiliary tasks at the same time. For example, in facial analysis, the primary task might be to predict facial landmarks, while the auxiliary tasks could be to estimate head poses and infer facial attributes Zhang et al. (2014). Since these tasks share common features, the model can use the training signals from the auxiliary tasks to improve its performance in the primary task.
Given that molecular properties are heavily influenced by molecular structure, it is reasonable to assume that representations useful for predicting one type of property could also help in predicting others. Based on this intuition, we propose to construct an auxiliary dataset of properties that can be computed using relatively inexpensive computational methods, but are not necessarily identical to the properties in the downstream datasets. Specifically, we select HOMO, LUMO, and Dipole Moment as the auxiliary properties because they can be accurately computed using Density Functional Theory (DFT). We also note that computing the auxiliary labels with the DFT is cheaper than obtaining more downstream labels via real-world experiments.
@@ 5 Experiments and Results
In this section, we address the following questions through a series of experiments: (1) Does the two-stage pretraining framework improve the downstream performance on datasets with limited labels? (2) How does each individual pretraining stage contribute to the improvements? (3) Is our assumpation that larger noise scales improve the generalization capability of the model correct? (4) Does the choice of pretraining dataset affect downstream performance? Additionally, we investigate how significant is the impact of finetuning dataset size to the downstream performance and the results are shown in the appendix.
@@ 5.1 Experiment Settings
We use GDB17 (Ruddigkeit et al., 2012) as the pretraining dataset for our model and other models to which we compare. We randomly select 1M unlabeled molecules from the 50M subset to be used in the first pretraining stage. We then sample 130K molecules out of the 1M subset to construct the auxiliary datasets for the second pretraining stage. The labels for the auxiliary dataset are computed with Psi4 (Smith et al., 2020). We use RDKit to generate 3D conformations from SMILES (Weininger, 1988) for models that take 3D graphs as inputs.
For benchmarking purposes, we use the same pretraining dataset to minimize any performance gains that might arise from the use of higher-quality pretraining datasets. Additionally, we use identical data splits for all pretraining methods to ensure fair and consistent comparisons.
5.2 Implementation Details
The primary and auxiliary encoders of MoleVers are built on the UniMol encoder architecture (Zhou et al., 2023). Each encoder comprises 15 layers, with an embedding dimension of 512 and a feedforward dimension of 2048. During the first pretraining stage, the model is trained for 1 million iterations using a batch size of 32, with a masking ratio of 0.15 for the MAP task. In the second pretraining stage, the model is trained for 50 epoch, maintaining the same batch size of 32. We employ the Adam optimizer with a learning rate of 10-4 and utilize a polynomial decay learning rate scheduler. We run all experiments on an NVIDIA Quadro RTX 8000 GPU.
Table 1: Quantitative results on the MPPW benchmark. We show the mean MAE ($\downarrow$) and R2 ($\uparrow$) across three train/test splits. Bolded and underlined values are the best and second best results, respectively. The numbers within the parentheses after the assay ids are the number of training molecules in the assay.
Assay GraphMVP MoleBert UniMol MolAE MoleVers
MAE R2 MAE R2 MAE R2 MAE R2 MAE R2
1 (50) 0.423 0.851 0.571 0.720 0.481 0.790 0.629 0.631 0.448 0.820
2 (50) 0.407 0.261 0.379 0.317 0.426 0.148 0.410 0.183 0.281 0.582
3 (47) 3.447 0.059 3.880 -0.175 3.393 -0.006 3.502 -0.048 3.103 0.192
4 (24) 0.329 0.716 0.591 0.098 0.475 0.292 0.590 0.075 0.319 0.727
5 (26) 0.463 -0.301 0.536 -0.772 0.440 -0.313 0.427 -0.259 0.417 -0.246
6 (48) 25.837 -0.520 23.255 -0.139 17.967 -0.018 18.190 -0.028 17.66 -0.012
7 (35) 0.655 0.168 0.945 -1.029 0.871 -0.800 0.763 -0.228 0.630 0.179
8 (30) 0.810 -0.829 0.728 -0.371 0.648 0.067 0.664 0.002 0.613 0.106
9 (34) 0.357 -0.092 0.413 -0.638 0.318 0.143 0.399 -0.264 0.312 0.170
10 (39) 0.233 0.307 0.258 0.159 0.211 0.387 0.271 0.092 0.187 0.490
11 (38) 0.426 0.121 0.671 -0.980 0.457 0.013 0.440 0.028 0.413 0.171
12 (25) 0.761 0.066 0.800 0.190 0.699 0.233 0.666 0.403 0.611 0.412
13 (22) 0.633 0.051 0.533 0.229 0.581 0.117 0.591 0.170 0.484 0.329
14 (43) 0.351 -0.240 0.267 0.377 0.331 0.099 0.303 0.217 0.280 0.314
15 (48) 0.397 0.656 0.477 0.564 0.524 0.439 0.495 0.473 0.385 0.665
16 (24) 0.885 0.279 0.855 0.287 0.910 -0.080 0.782 0.301 0.700 0.364
17 (42) 1.164 0.170 1.315 -0.593 1.277 -0.098 1.385 -0.256 1.142 0.143
18 (31) 0.195 0.693 0.170 0.732 0.202 0.685 0.202 0.617 0.141 0.855
19 (62) 0.353 -1.002 0.265 -0.330 0.258 -0.193 0.230 -0.084 0.191 0.259
20 (51) 0.347 -0.557 0.259 0.260 0.294 0.010 0.330 -0.217 0.234 0.343
21 (19) 0.361 0.570 0.507 0.224 0.620 -0.354 0.493 0.155 0.351 0.572
22 (22) 0.983 -0.335 0.733 0.201 0.608 0.241 0.580 0.266 0.526 0.263
@@ 5.3 Results on the MPPW Benchmark
In the MPPW benchmark, we compare MoleVers with four baselines: state-of-the-art GNNs, GraphMVP (Liu et al., 2022) and Mole-BERT (Xia et al., 2023), as well as state-of-the-art transformers, Uni-Mol (Zhou et al., 2023) and Mol-AE (Yang et al., 2024). All models are implemented in PyTorch (Paszke et al., 2019) and trained from scratch using publicly available source code. We also provide comparisons with more baselines on large downstream datasets in Section 5.7.
For each downstream dataset, we construct three distinct train/test splits with a 1:1 train-test ratio. All models are finetuned for 50 epochs on the training splits, and the downstream performance of the last epoch is recorded in Table 1. We evaluate the downstream performance using two metrics: mean absolute error (MAE) and the coefficient of determination (R2), which indicate how well the model can explain the variance in the data.
As shown in Table 1, the R2 scores for the current state-of-the-art models are relatively low. This indicates that existing models could not consistently learn molecular representations that are useful for property prediction. This highlights the need for more effective pretraining methods suited to low-data regimes. In contrast, MoleVers outperforms other baseline models in 20 out of the 22 assays, and achieving a close second rank in the remaining two. Notably, no other method consistently ranks among the top two across all assays. These results demonstrate that the two-stage pretraining framework is an effective approach for improving downstream performance when labeled data is extremely limited.
One might argue that the performance gains shown in Table 1 are simply due to the additional labels, as other methods are pretrained without auxiliary labels. To address this concern, we have conducted further experiments, detailed in Table 8 in the Appendix, where other models are also pretrained with the auxiliary labels. In such a setup too, our model achieves state-of-the-art MAE in 20 out of 22 assays, while placing a close second in the remaining two. This demonstrates that the gains achieved by MoleVers are not solely because of the additional data, but rather stem from the synergy between the two pretraining stages. Additionally, we emphasize that the second pretraining stage offers a cost-effective solution for improving downstream performance, as the computational cost of obtaining auxiliary labels is negligible compared to the costs of acquiring downstream labels through wet-lab experiments.
@@ 5.4 Ablation of Pretraining Stages
We study the influence of each pretraining stage on the downstream performance of MoleVers through a series of ablation studies. As shown in Table 2, incorporating either the first or second pretraining stage into the pipeline always leads to better downstream performance compared with directly training the model on the downstream datasets. Interestingly, the improvements vary across assays: some benefit more from the first pretraining stage, while others see more gains from the second pretraining stage. This variation could be due to the auxiliary properties we have chosen–HOMO, LUMO, and Dipole Moment–which are more related to intrinsic molecular properties (e.g., assay 1), rather than complex interactions (e.g., assay 3). Overall, the combination of both pretraining stages consistently yields the best downstream performance across all assays.
Table 2: Ablation studies of our pretraining strategy. We can see that combining both pretraining stage 1 and stage 2 gives the best performance on the downstream datasets.
@@
MAE R2 MAE R2 MAE R2 MAE R2
- - 0.683 0.595 0.493 0.032 3.680 -0.063 0.784 -0.700
- 0.592 0.680 0.420 0.192 3.161 0.086 0.431 0.479
- 0.501 0.771 0.343 0.418 3.301 0.081 0.346 0.635
0.448 0.820 0.281 0.582 3.103 0.192 0.319 0.727
Table 3: Effects of noise scales on downstream performance. Larger noise scales tend to improve the downstream performance of MoleVers. However, using excessively large noise scales (e.g., max. $\sigma$ = 20) leads to training instability.
@@ 5.5 Impact of Noise Scale on Downstream Performance
In Section 3.1.2, we hypothesized that using larger noise scales for the denoising tasks can improve the downstream performance. In Table 3, we show the downstream performance of MoleVers with various noise scales drawn from a uniform distribution, $\sigma\sim\mathcal{U}(0,b)$, where $b$ is the maximum noise scale. Note that, similar to what has been observed in a prior work (Yang et al., 2024), the pretraining become unstable when excessively larger noise scales, e.g., $b=20$, are used. Therefore, we limit our ablation study to a maximum value of 10.
@@ 5.6 Impact of Pretraining Dataset Quality on Downstream Performance
Table 4: Impact of pretraining (stage 1) dataset diversity, measured by the number of training samples. The downstream performance of MoleVers improves as the number of training samples increases.
Dataset size Assay ID
(number of 1 2 3 4
training samples) MAE R2 MAE R2 MAE R2 MAE R2
10,000 1.152 -0.008 0.498 -0.083 3.660 -0.047 0.611 -0.016
100,000 0.629 0.636 0.409 0.198 3.205 0.103 0.549 0.189
1,000,000 0.592 0.680 0.420 0.192 3.161 0.086 0.431 0.479
Table 5: Impacts of pretraining (stage 1) dataset diversity, measured by the variety of atom types. The downstream performance of MoleVers improves when the number of unique atom types in the training set increases.
@@
C N O F Misc. MAE R2 MAE R2 MAE R2 MAE R2
- - - 1.093 0.018 0.496 -0.052 3.584 -0.013 0.628 -0.015
- - 0.845 0.426 0.431 0.191 3.428 0.041 0.480 0.416
- 0.619 0.601 0.423 0.233 3.273 0.055 0.493 0.281
0.592 0.680 0.420 0.192 3.161 0.086 0.431 0.479
5.7 Results on Large Downstream Datasets
We further evaluate the performance of MoleVers on the MoleculeNet benchmark (Wu et al., 2018), focusing on large-scale regression datasets such as QM7, QM8, and QM9. These datasets, which range from thousands to over a hundred thousand labeled molecules, provide insights into the effectiveness of our pretraining strategy in data-abundant scenarios. As shown in Table 6, MoleVers outperforms all baseline models across all datasets, achieving the lowest MAE scores. Therefore, the proposed two-stage pretraining framework is not only effective in low-data regimes, but also excels when abundant labeled data is available.
. Dataset QM7 QM8 QM9 #Molecules 6830 21789 133885 D-MPNN 103.5 0.0190 0.0081 Attentive FP 72.0 0.0179 0.0081 GROVER 94.5 0.0218 0.0099 MolCLR 66.8 0.0178 - Uni-Mol 58.9 0.0160 0.0054 Mol-AE 53.8 0.0161 0.0053 MoleVers (ours) 51.3 0.0155 0.0050
@@ 6 Conclusion
In this work, we addressed the challenge of molecular property prediction in the wild, i.e., in real-world scenarios where molecular property labels that are validated through experiments are scarce. We introduced a two-stage pretraining strategy that employs masked atom prediction, dynamic denoising, and auxiliary property prediction to learn robust molecular representations. To enable effective denoising pretraining with larger noise scales, we proposed a novel branching encoder that decouples the MAP pipeline from the denoising pipeline. We evaluated our model on a new benchmark, Molecular Property Prediction in the Wild, designed to reflect real-world data limitations. Our model consistently outperforms previous state-of-the-art baselines in both low-data and high-data regimes. Our results highlight the effectiveness of the two-stage pretraining strategy, making it suitable for real-world applications where labeled data are extremely limited.

### v2

@@ 3.2 Stage 2: Auxiliary Property Prediction
We further improve the generalization capability of the primary encoder by incorporating auxiliary property prediction in the second pretraining stage. This approach is inspired by multi-task learning (Caruana, 1997), where a model is trained to solve both the primary task and related auxiliary tasks at the same time. For example, in facial analysis, the primary task might be to predict facial landmarks, while the auxiliary tasks could be to estimate head poses and infer facial attributes (Zhang et al., 2014). Since these tasks share common features, the model can use the training signals from the auxiliary tasks to improve its performance in the primary task.
Given that molecular properties are heavily influenced by molecular structure, it is reasonable to assume that representations useful for predicting one type of property could also help in predicting others. Based on this intuition, we propose to construct an auxiliary dataset of properties that can be computed using relatively inexpensive computational methods, but are not necessarily identical to the properties in the downstream datasets. Specifically, we select highest occupied molecular orbital (HOMO), lowest unoccupied molecular orbital (LUMO), and dipole moment as the auxiliary properties because they can be accurately computed using density functional theory (DFT). We also note that computing the auxiliary labels with DFT is significantly cheaper than obtaining more downstream labels via real-world experiments.
@@
3.2.1 Auxiliary Label Generation with LLM
Another promising approach in our second pretraining stage is to leverage large language models (LLMs) to generate the auxiliary labels. LLMs have demonstrated strong generalization capability across a wide range of tasks due to their exposure to diverse pretraining data. As a result, LLMs can be used to generate molecular property labels directly within the downstream task space with a fraction of computational cost of DFT.
Directly predicting the absolute values of properties using LLMs, however, often results in lower-quality labels. Instead of absolute values, we propose to generate relative-valued labels in the form of pairwise rankings of molecules. Pairwise rankings reformulate the problem as a binary classification task, where the input consists of two molecules and the output is either 0 or 1, indicating which molecule has a higher value for the given property. These outputs can then be used either as auxiliary labels in the second pretraining stage, or as additional targets during finetuning (Figure 2). We provide a more thorough discussion of LLM-based label generation in Appendix A.1.
@@ 5 Experiments and Results
In this section, we address the following questions through a series of experiments: (1) Does the two-stage pretraining framework, including LLM-based generated labels, improve the downstream performance on datasets with significantly limited labels? (2) How does each individual pretraining stage contribute to the improvements? (3) Is our assumption that larger noise scales improve the generalization capability of the model correct? (4) Does the choice of pretraining dataset affect downstream performance? Additionally, we investigate how significant the impact of finetuning dataset size is to the downstream performance, for which the results are shown in the appendix.
Figure 3: Box plot of (a) mean absolute error ($\downarrow$) and (b) R2 score ($\uparrow$) on the 22 assays and 3 train/test splits in the MPPW benchmark. The whiskers extend from the box to the farthest data point lying within 1.5x the inter-quartile range from the box. We can see that MoleVers variants, especially MoleVers, achieve the best overall performance with significant margins in both metrics. The full results are shown in Table 9 in Appendix.
Figure 4: Number of assays in the MPPW benchmark (22 assays, 3 train/test splits) in which a model achieves the best or second best performance in terms of mean absolute error. Molevers demonstrates state-of-the-art performance, outperforming the baselines in 18 assays and achieving the second-best performance in the remaining 4 assays.
@@ 5.1 Experiment Settings
Datasets. We use GDB17 (Ruddigkeit et al., 2012) as the pretraining dataset for our model and other baselines to minimize any performance differences that might arise from the use of higher-quality pretraining datasets. We randomly select 1M unlabeled molecules from the 50M subset to be used in the first pretraining stage. We then sample 130K molecules out of the 1M subset to construct the auxiliary datasets for the second pretraining stage. The labels for the DFT-based auxiliary dataset are computed with Psi4 (Smith et al., 2020), while ChatGPT4o is used for the LLM-based auxiliary dataset. We use RDKit to generate 3D conformations from SMILES (Weininger, 1988) for models that take 3D graphs as input.
Models & baselines. We provide evaluation results of four baselines: state-of-the-art GNNs, GraphMVP (Liu et al., 2022) and Mole-BERT (Xia et al., 2023), as well as state-of-the-art transformers, Uni-Mol (Zhou et al., 2023) and Mol-AE (Yang et al., 2024). All models are implemented in PyTorch (Paszke et al., 2019) and trained from scratch using publicly available source code. Afterwards, the models are finetuned for 50 epochs on the downstream train split, and the downstream test split performance of the last epoch is recorded in Table 9. There are three train/test splits for each downstream dataset in the benchmark with a 1:1 train-test ratio. We also provide comparisons with more baselines on large downstream datasets in Appendix A.5. More details about network architecture and training strategy are available in Appendix A.2.
We evaluated four variants of our proposed model, MoleVers, on the proposed MPPW benchmark. These four variants are different in second pretraining stage and finetuning strategies: (1) MoleVersD2, pretrained on 130K labels generated by density functional theory (DFT) during the second stage, (2) MoleVersL2, pretrained on 1M pairs of molecule ranking with labels generated by a large language model (LLM) during the second stage, (3) MoleVersD2L3, pretrained on 130K DFT-generated labels during the second stage, then additionally finetuned using LLM-based ranking labels of the downstream test-split molecules, (4) MoleVers, pretrained on 130K DFT-generated labels. For this latest variant, LLM-generated ranking labels are only used if the absolute value of the rank correlation of the LLM predictions on the downstream training split is greater than 0.4. Note that we finetune all variants on downstream train labels.
Evaluation metrics. We use two metrics for evaluation: mean absolute error (MAE) and the coefficient of determination (R2). Models with perfect predictions would achieve R2 scores of 1, while simply predicting the mean of the labels would result in an R2 score of 0.
@@ 5.2 Results on the MPPW Benchmark
The predictive performance of the baseline models, evaluated on the MPPW benchmark, is presented in Figure 3 (and extensively in Table 9 in Appendix). From Figure 3, the poor R2 scores of the baselines, all below 0.14, indicate that their predictions are only marginally better than simply predicting the mean values of the properties. This highlights the need for more effective pretraining methods suited to small-data regimes. Note that Mole-BERT performs poorly on a few datasets, resulting in a significantly low average R2 score.
All variants of MoleVers outperform the baselines in terms of both average MAE and R2 scores. Figure 4 shows that MoleVers L2, pretrained exclusively with LLM-generated auxiliary labels, surpasses all baselines in 10 out of the 22 assays, and ranks second in 5 of the remaining assays. It also achieves improvements of 8.42% in average MAE and 97.84% in average R2 score compared to the best-performing baseline. This proves that a general-purpose LLM can be leveraged to generate weak labels for molecular property predictions. Such an approach can be useful in scenarios where access to more accurate tools, such as DFT, is limited due to resource or time constraints.
However, the availability of LLM-generated auxiliary labels does not always yield the optimal performance, particularly if more accurate auxiliary labels are accessible. For example, MoleVers D2, pretrained using DFT-based auxiliary labels, outperforms MoleVers D2L3, which incorporates LLM-based labels during finetuning. Nonetheless, MoleVers D2L3 still surpasses all baselines, outperforming them in 11 out of the 22 assays and ranking second in 8 of the remaining assays.
Interestingly, we observed that incorporating LLM-based labels during finetuning can improve performance in some assays. We hypothesize that the improvement correlates with the quality of the LLM predictions for a given assay. To test this hypothesis, we incorporate LLM-based labels during finetuning only when the absolute rank correlation coefficient of the LLM predictions on the training set is above 0.4. This strategy results in our main model, MoleVers, which outperforms the baselines in 18 out of the 22 assays and ranks second in the remaining 4 assays. MoleVers demonstrates improvements of 13.67% in average MAE and 133.09% in average R2 score compared to the best-performing baseline. Overall, these results confirm that the two-stage pretraining framework is an effective approach for improving downstream performance when labeled data is extremely limited.
@@ 5.3 Ablation of Pretraining Stages
We study the influence of each pretraining stage on the downstream performance of MoleVers through a series of ablation studies. We use MoleVers D2 in all of our ablation studies. As shown in Table 1, incorporating either the first or second pretraining stage into the pipeline always leads to better downstream performance compared with directly training the model on the downstream datasets. Interestingly, the improvements vary across assays: some benefit more from the first pretraining stage, while others see more gains from the second pretraining stage. This variation could be due to the auxiliary properties we have chosen–HOMO, LUMO, and Dipole Moment–which are more related to intrinsic molecular properties (e.g., assay 1), rather than complex interactions (e.g., assay 3). Overall, the combination of both pretraining stages consistently yields the best downstream performance across all assays.
Table 1: Ablation studies of our pretraining strategy. We report the mean MAE ($\downarrow$) across three train/test splits. We can see that combining both pretraining stage 1 and stage 2 gives the best performance on the downstream datasets.
@@
- - 0.683 0.493 3.680 0.784
- 0.592 0.420 3.161 0.431
- 0.501 0.343 3.301 0.346
0.417 0.298 2.999 0.337
5.4 Ablation of Branching Encoder and Extreme Denoising
The key components that enable denoising pretraining with higher noise levels in the first stage are the branching encoder and dynamic noise scale sampling. Here, we study the impact of each component to the downstream performance. As shown in Table 2, using a single encoder for denoising pretraining at larger noise levels generally leads to worse prediction performance. In contrast, the introduction of the branching encoder can mitigate this issue in most cases. Furthermore, combining the branching encoder with dynamic noise scale sampling consistently yields the best downstream performace, highlighting the importance of dynamic over static noise scales.
Table 2: Ablation studies of the proposed branching encoder and extreme denoising. B.E. and D.D. stands for branching encoder and dynamic denoising, respectively. We report the mean MAE ($\downarrow$) across 3 train/test splits. Combining branching encoder with varying noise levels of a broader range yields the best downstream performance.
B.E. D.D. Max $\sigma$ Assay ID
1 2 3 4
- - 1 0.481 0.426 3.393 0.475
- - 10 0.519 0.418 3.401 0.492
- 10 0.521 0.336 3.301 0.476
10 0.417 0.298 2.999 0.337
Table 3: Effects of noise scales on downstream performance. We report the mean MAE ($\downarrow$) across three train/test splits. Larger noise scales tend to improve the downstream performance of MoleVers. However, using excessively large noise scales (e.g., max. $\sigma$ = 20) leads to unstable training.
@@ 5.5 Impact of Noise Scale on Downstream Performance
In Section 3.1.2, we hypothesized that using larger noise scales for the denoising tasks can improve the downstream performance. In Table 3, we show the downstream performance of MoleVers with various noise scales.Note that, similar to what has been observed in a prior work (Yang et al., 2024), the pretraining become unstable when excessively larger noise scales, e.g., $b=20$, are used. Therefore, we limit our ablation to a maximum value of 10.
@@ 5.6 Impact of Pretraining Dataset Quality on Downstream Performance
Table 4: Impact of dataset diversity (stage 1), measured by the number of training samples. We report the mean MAE ($\downarrow$) across 3 train/test splits. The performance of MoleVers improves as the number of training samples increases.
Training size Assay ID
1 2 3 4
10,000 1.152 0.498 3.660 0.611
100,000 0.629 0.409 3.205 0.549
1,000,000 0.592 0.420 3.161 0.431
Table 5: Impacts of pretraining (stage 1) dataset diversity, measured by the variety of atom types. The number of molecules in each dataset is fixed to 100K for a fair comparison. We report the mean MAE ($\downarrow$) across 3 train/test splits. M stands for miscellaneous atom types. The downstream performance of MoleVers improves when the number of unique atom types in the training set increases.
@@
C N O F M.
- - - 1.093 0.496 3.584 0.628
- - 0.845 0.431 3.428 0.480
- 0.619 0.423 3.273 0.493
0.592 0.420 3.161 0.431
@@ 6 Conclusion
We addressed the challenge of molecular property prediction in the wild—scenarios where experimentally-validated labels are significantly limited—using a two-stage pretraining strategy. In the first stage, we introduced a novel extreme denoising pretraining task, enabled by our branching encoder and dynamic noise scale sampling. In the second stage, we pretrained the model with auxiliary labels obtained through high-accuracy methods such as density functional theory, or more computationally efficient approaches such as large language models. We evaluated our model on a new benchmark, Molecular Property Prediction in the Wild, with extremely small datasets that were chosen to reflect real-world data limitations. Our model consistently outperforms existing baselines, making it suitable for real-world applications where labeled data are limited.